OpenStudy (anonymous):

Out of all these functions which ones are even? y = sin x y = csc x y = sec x y = cos x

OpenStudy (anonymous):

even function : f(-x) = f(x) can you do it now ?

OpenStudy (anonymous):

y= cos x y= sec x

OpenStudy (anonymous):

@prawal don't give answers

OpenStudy (anonymous):

Cool secto... when you say f(-x) = f(x) ..what do you mean?

OpenStudy (anonymous):

for example x^2 is an even function since f(-x) = (-x)^2 = x^2 = f(x)

OpenStudy (anonymous):

how can f(-x) = f(x)?

OpenStudy (anonymous):

@Zekarias @Coolsector already gave the explanation.. i do no see any need to do it again...

OpenStudy (anonymous):

if f(-x) = f(x) then f(x) is an even function i just gave you an example which f(x) = f(-x) so it is possible

OpenStudy (anonymous):

OH sorryI misunderstood the explanation ..ty

OpenStudy (anonymous):

even function are symmetric with respect to the y axis

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw .. :)